Equation of the ellipse whose minor axis is equal to the distance between foci and whose latus rectum is 10, is given byA) 2x2+3y2=100B) 2x2+3y2=80C) x2+2y2=100D) none of these

Grade:

1 Answers

Arun
6 years ago
Dear student
 
2b = 2ae
b = ae
 
and 
 
2 b^2/ a = 10
b^2 = 5a
 
when solving the equation
 
b^2 = 50 and a^2 = 100
 
now 
equation of ellipse
 
x^2 / 100  +  y^2 / 50 = 1
 
 
Regards
Arun (askIITians forum expert)